Here are summary statistics for randomly selected weights of newborn​ girls:

nequals=170

x overbarxequals=29.6

​hg,

sequals=7.1

hg. Construct a confidence interval estimate of the mean. Use a

99%

confidence level. Are these results very different from the confidence interval

27.6

hgless than<muμless than<31.2

hg with only

19

sample​ values,

x overbarxequals=29.4

​hg, and

sequals=2.7

​hg?

What is the confidence interval for the population mean

muμ​?

hgless than<muμless than<

hg ​(Round to one decimal place as​ needed.)

Are the results between the two confidence intervals very​ different?

A.

​No, because the confidence interval limits are similar.

B.

​No, because each confidence interval contains the mean of the other confidence interval.

C.

​Yes, because one confidence interval does not contain the mean of the other confidence interval.

D.

​Yes, because the confidence interval limits are not similar.
Answer #1

0 - no 170 X = 29,6 hg S 7.1 hg 99% confidence interval estimate of the mean is, PI X - Zarz S s els X + Z012 s vo 15 = 0.99

n=19, x=29,4 2) S - 2.7 Z2U/2 = 20.005 = 2,51 X + Zala S Upper limite - 29,4 + 2,57 x 2.7 Nig 29.44 l. 5919 - 30.9919 Lowex l